# A 26-Gram Butterfly-Inspired Robot Achieving Autonomous Tailless Flight

### v1

@@ Abstract
Flapping-wing micro air vehicles (FWMAVs) have demonstrated remarkable bio-inspired agility, yet tailless two-winged configurations remain largely unexplored due to their complex fluid-structure and wing-body coupling. Here we present AirPulse, a 26-gram butterfly-inspired FWMAV that achieves fully onboard, closed-loop, untethered flight without auxiliary control surfaces. The AirPulse robot replicates key biomechanical traits of butterfly flight, including low wing aspect ratio, compliant carbon-fiber-reinforced wings, and low-frequency, high-amplitude flapping that induces cyclic variations in the center of gravity and moment of inertia, producing characteristic body undulation. We establish a quantitative mapping between flapping modulation parameters and force-torque generation, and introduce the Stroke Timing Asymmetry Rhythm (STAR) generator, enabling smooth, stable, and linearly parameterized wingstroke asymmetry for flapping control. Integrating these with an attitude controller, the AirPulse robot maintains pitch and yaw stability despite strong oscillatory dynamics. Free-flight experiments demonstrate stable climbing and turning maneuvers via either angle offset or stroke timing modulation, marking the first onboard controlled flight of the lightest two-winged, tailless butterfly-inspired FWMAV reported in peer-reviewed literature. This work corroborates a foundational platform for lightweight, collision-proof FWMAVs, bridging biological inspiration with practical aerial robotics. Their non-invasive maneuverability is ideally suited for real-world applications, such as confined-space inspection and ecological monitoring, inaccessible to traditional drones, while their biomechanical fidelity provides a physical model to decode the principles underlying the erratic yet efficient flight of real butterflies.
@@ Introduction
Flapping-wing Micro Aerial Vehicles have long drawn inspiration from nature with the aim of replicating the agility, efficiency, and maneuverability of flying animals  (?). These robots are particularly favorable for small-scale aerial platforms, as rotary- and fixed-wing counterparts suffer from strong Reynolds effects at low scales, whereas flapping wings retain aerodynamic efficiency through unsteady lift mechanisms. Considerable progress has been achieved in insect-scale designs using high-bandwidth soft actuators such as piezoelectric bimorphs  (?) and dielectric elastomer actuators  (?), enabling demonstrations of robotic fruit flies and bees  (?, ?, ?, ?, ?, ?). However, these sub-gram platforms remain constrained by limited flight duration, offboard power reliance, and minimal payload capacity, mainly due to the limitations rooted in the absence of engineered materials and fabrication methods capable of replicating insect biomechanics. To overcome this, researchers have designed small-scale FWMAVs below 100 g  (?), leveraging their larger size to integrate onboard computation and power while maintaining strict Size, Weight, and Power (SWaP) constraints. These systems, inspired by birds  (?, ?, ?, ?, ?, ?, ?), bats  (?, ?, ?, ?), beetles  (?, ?, ?), dragonflies  (?), fruit flies and more  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?), have extended operational endurance and flight complexity, marking an important evolution in flapping-wing robotics.
@@
Amidst these advances, one of nature’s most aerodynamically sophisticated fliers, the butterfly, remains surprisingly underexplored. Butterflies exhibit a unique flight paradigm that is distinct from that of fast-flapping insects and larger vertebrate flyers. Unlike birds and bats, which stabilize and maneuver through active wing morphing or tail actuation  (?, ?, ?), butterflies achieve controlled tailless flight using highly compliant wings driven anteromotorically by forewings. In contrast to bees and fruit flies that flap above 150 Hz  (?), butterflies generate sufficient aerodynamic force with large-amplitude strokes at only $\sim$10 Hz  (?, ?). Their low wing aspect ratios (typically <5<5) yield the largest wing area-to-body-mass ratios among flying animals  (?). Despite these seemingly unfavorable traits, butterflies show remarkable agility—executing rapid climbs, sharp turns, and erratic zigzags  (?)—demonstrating a flight strategy that prioritizes stability through dynamic coupling rather than high flapping rate or auxiliary control surfaces.
@@
Existing aerodynamic studies have focused primarily on insects with rigid, high-aspect-ratio wings flapping at high frequencies. Although butterflies also exploit unsteady mechanisms such as delayed stall  (?), wake capture  (?), and clap-and-fling  (?), their biomechanical realization of these effects differs fundamentally. Their large and compliant wings interact strongly with airflow, leading to complex fluid-structure coupling and increased thoracic loading that ultimately lowers flapping frequency while amplifying unsteady effects. Despite the vast diversity of wing morphologies  (?), the aerodynamic contributions of these traits remain poorly characterized  (?). The anteromotoric flight of butterflies further renders them inherently underactuated in six-degree-of-freedom (6-DOF), posing a long-standing challenge in understanding how tailless animals achieve maneuvers such as turning  (?). Moreover, their characteristic body undulation likely arises from intricate wing-body interactions that couple aerodynamic and inertial effects  (?, ?, ?).
@@
While biological investigations using high-speed videography have revealed complex flight behaviors such as climbing  (?, ?), forward flight  (?, ?), and turning  (?, ?), these studies remain descriptive and lack quantitative links to controllable actuation. On the other side, modern Computational Fluid Dynamics (CFD) simulations can provide insight into flow structures  (?, ?), but they often assume rigid wings and omit aeroelastic coupling and body-wing interaction, limiting their biological and engineering relevance. Due to the rudimentary understanding of biomechanics and the difficulties of modeling fluid-structure interactions and wing-body coupling  (?, ?, ?), the translation of the unique mechanisms underlying butterfly flight into engineered systems has remained limited  (?, ?, ?, ?, ?). Butterfly-inspired, two-winged, tailless FWMAVs—especially those with low-aspect-ratio wings and low flapping frequencies—remain largely unexplored in robotics. This knowledge gap presents both a challenge and an opportunity to uncover the principles governing butterfly flight and translate them into controllable flapping-wing robots.
@@
In this work, we present AirPulse, a 26 g butterfly-inspired flapping-wing robot that, to the best of our knowledge, represents the first and lightest two-winged tailless FWMAV with fully onboard sensing and closed-loop control reported in peer-reviewed work (Fig. 1, A and B, Table 1, and movie S1). Despite the abundance of research on four-winged and tailed flapping-wing platforms, two-winged tailless systems remain poorly understood, primarily because their flight dynamics are dominated by strong fluid-structure coupling and the absence of tail-based stabilization. The AirPulse robot addresses this gap by replicating the fundamental aerodynamic and structural characteristics of butterflies, featuring low wing aspect ratio, biologically relevant wing loading—the lowest among reported studies, indicating efficient aerodynamics (Table 1)—and highly compliant wings reinforced by carbon fiber spars in a vein-inspired layout (Fig. 1G). This biomimetic reinforcement yields spatially heterogeneous stiffness, allowing compliant yet controllable wing deformation, thereby contrasting the uniform stiffness typical of rod-reinforced or rigid-wing designs  (?, ?). We also revealed that the large-amplitude flapping motion induces cyclic variations in the center of gravity and moment of inertia (Fig. 2, A and B), which manifest as pronounced body undulation, a critical challenge for control in tailless flapping-wing flight.
@@
A central obstacle in controlling such underactuated systems lies in the lack of a quantitative relationship between flapping modulation and force-torque generation, a gap that prior studies have not explicitly addressed  (?, ?, ?, ?, ?). We systematically characterized the mapping between flapping wing modulation parameters—including amplitude, frequency, stroke angle offset, and stroke timing—and the resulting forces and torques (Fig. 4). This mapping fills a fundamental gap in understanding how such tailless flapping-wing robots allocate control authority in the absence of auxiliary surfaces, establishing a direct link between modulation strategy and maneuvering capability. Furthermore, we introduced a novel rhythm generator, namely Stroke Timing Asymmetry Rhythm (STAR), which enables smooth and stable asymmetric wing motion, where the degree of asymmetry varies linearly with a single control parameter (Fig. 3, E and F). This formulation directly overcomes the discontinuity and instability issues found in prior stroke timing modulation approaches  (?, ?, ?, ?), providing a mathematically grounded and physically robust foundation for precise stroke timing control.
@@
Building upon these insights, we developed an onboard control architecture that integrates Proportional-Integral-Derivative (PID)-based attitude regulation with real-time state estimation to cope with the inherent oscillatory dynamics of undulatory flight. In contrast to fixed-wing or bird-inspired robots that rely on steady aerodynamic moments, the AirPulse robot must continuously manage rapidly varying inertial and aerodynamic couplings throughout each wingbeat. Our controller was shown to maintain pitch and yaw stability even under large oscillatory motion of the body. Experimental validation demonstrated fully untethered, stable free flight using two flapping modulations (angle offset and stroke timing based on STAR)—without tails or external control surfaces—achieving pitch and yaw tracking, climbing, and turning maneuvers with an average power consumption of $\sim$5.9 W and a power loading of 4.38 g/W (Fig. 5, D and E, and movie S4 and S5). These results showcase the first onboard controlled flight of a butterfly-inspired robot of 26 g and establish a foundation for understanding and controlling tailless, underactuated flapping-wing systems with two low aspect ratio wings, thus paving the way for lightweight and power-efficient micro aerial vehicles suited for confined environments such as greenhouse inspection and ecological monitoring.
@@ Results
To investigate how butterfly-like wing kinematics and flapping modulation enable controlled free flight, we developed AirPulse (Fig. 1A), a small-scale flapping-wing robot inspired by the lime swallowtail (Papilio demoleus), one of the most widely distributed butterfly species worldwide. To the best of our knowledge, AirPulse is the first and lightest butterfly-inspired flapping-wing platform reported to date with fully integrated onboard sensing and closed-loop control (flight-ready mass $\sim$26 g; dry mass $\sim$24 g; Fig. 1B and Table 1). The robot was designed as a biological counterpart, being lightweight, morphologically plausible, and instrumented for quantitative studies of aeroelastic wing-air interactions, flapping modulation, and feedback control.
@@
The AirPulse robot features a biomimetic wing structure (wingspan $\sim$60 mm) connected to a compact body core (Fig. 1A and movie S1). Its lightweight design was achieved primarily through two complementary strategies. First, we adopted a novel body architecture in which all avionics are directly assembled on a single 3D-printed fuselage (Fig. 1C), unlike previous designs where the control unit or battery was mounted at the abdomen and connected to the thorax via a carbon rod  (?, ?, ?, ?). This integration eliminates structural redundancy and allows the Inertial Measurement Unit (IMU) to be co-located near the Center of Gravity (CG). Second, we minimized the wing surface area, yielding the smallest wing planform reported among butterfly-inspired robots in the literature (Table 1). Guided by scaling laws and geometric similarity  (?), where $S\sim l^{2}$ and $m\sim l^{3}$ ($S$ is single wing area, $m$ is body mass, $l$ is characteristic length), we estimated the desired wing size based on the biologically observed mass ratio between wings and body. Specifically, we calculated the non-wing mass and divided it by 0.7, assuming the wings contribute approximately 30% of total mass in butterflies  (?). This provided an estimate of the total mass and required wing surface. Starting from a larger prototype capable of generating sufficient lift, we scaled the design accordingly and iteratively refined it through flight tests to identify the minimal wing area that maintained stable lift and control. This iterative process substantially reduced the overall mass while preserving flight capability. As a result, the wings of the AirPulse robot account for approximately 38% of the total body mass (Fig. 1B), comparable to that observed in dried specimens of Papilio xuthus  (?), confirming the biological plausibility of the design.
@@
Each lateral wing assembly comprises a forewing and hindwing mechanically coupled via a 3D-printed connector to a common servo arm. Two micro servos mounted on the body core independently actuate the forewings (Fig. 1D), reproducing the anteromotoric wing actuation seen in real butterflies  (?). This configuration yields a coupled, synchronous forewing-hindwing motion with a characteristic phase lag (movie S2). For efficient actuation, the servo axes are canted relative to the longitudinal symmetry plane to align the torque transmission with the structural stiff axis. This configuration produces a fixed $50^{\circ}$ V-shaped dihedral, realized through the fuselage and servo mounts, which promotes inter-wing force asymmetry similar to that of real butterflies  (?). The resulting flapping motion passively induces body undulation with a phase offset between flapping and body pitch angle (Fig. 1D and Fig. 5, F and H). Due to the larger wing-to-body ratio of the robot, this oscillatory coupling is pronounced than in larger robotic fliers (bird-inspired drones), motivating control strategies that explicitly account for periodic body dynamics as discussed later.
The body core houses a custom flight control board featuring a 9-axis IMU and barometer for state estimation, direct Pulse Width Modulation (PWM) servo control, and per-rail current monitoring (Fig. 1C and Materials and Methods). All states are telemetered to the ground station for data logging. The IMU is positioned with the robot’s CG and aligned to the body frame (+x+x forward, +y+y left, +z+z up) to minimize inertial coupling and aerodynamic interference. A lightweight power management circuit supports both 1S and 2S LiPo batteries, allowing weight reduction with a trade-off in flight endurance.
@@
For structural layout, we analyzed the venation morphology across 24 specimens from three major families: Nymphalidae, Pieridae, and Papillionidae (Materials and Methods). Forewing venation typically consists of a discal cell (Dc) and five sets of longitudinal veins—subcostal (Sc), radial (R), medial (M), cubital (Cu), and anal (A)—following the Comstock-Needham system  (?) (Fig. 1E). Among families, we found that the Dc-Cu1 angle varied systematically (Fig. 1F and fig. S2): Pieridae showed the largest mean (131.3° ± 3.5°), followed by Nymphalidae (124.3° ± 4.6°) and Papillionidae (123.2° ± 5.8°). Aspect ratio was highest in Papillionidae (0.85 ± 0.14) and lowest in Pieridae (0.71 ± 0.05). Regression analysis revealed a negative correlation between Dc-Cu1 angle and the reciprocal of aspect ratio across all species (slope = –0.0117, $R^{2}$ = 0.40, $p$ = 0.0009), suggesting that larger Dc-Cu1 angles correspond to higher aspect ratios. These morphological couplings indicate that vein geometry influences aerodynamic shape and load-bearing characteristics, serving as principles directly transferrable to biomimetic wing design.

### v2

@@ Abstract
The flight of biological butterflies represents a unique aerodynamic regime where high-amplitude, low-frequency wingstrokes induce significant body undulations and inertial fluctuations. While existing tailless flapping-wing micro air vehicles typically employ high-frequency kinematics to minimize such perturbations, the lepidopteran flight envelope remains a challenging and underexplored frontier for autonomous robotics. Here, we present AirPulse, a 26-gram butterfly-inspired robot that achieves the first onboard, closed-loop controlled flight for a tailless two-winged platform at this scale. It replicates key biomechanical traits of butterfly flight, utilizing low-aspect-ratio, compliant carbon-fiber-reinforced wings and low-frequency flapping that reproduces characteristic biological body undulations. Leveraging a quantitative mapping of control effectiveness, we introduce a hierarchical control architecture featuring state estimator, attitude controller, and central pattern generator with Stroke Timing Asymmetry Rhythm (STAR), which translates attitude control demands into smooth and stable wingstroke timing and angle-offset modulations. Free-flight experiments demonstrate stable climbing and directed turning maneuvers, proving that autonomous locomotion is achievable even within oscillatory dynamical regimes. By bridging biological morphology with a minimalist control architecture, AirPulse serves as both a hardware-validated model for decoding butterfly flight dynamics and a prototype for a new class of collision-resilient aerial robots. Its lightweight and compliant structure offers a non-invasive solution for a wide range of applications, such as ecological monitoring and confined-space inspection, where traditional drones may fall short.
@@ Introduction
Flapping-wing Micro Aerial Vehicles have long drawn inspiration from nature to replicate the agility, efficiency, and maneuverability of flying animals [1]. At micro-scales, flapping wings are particularly advantageous as they leverage unsteady lift mechanisms to maintain aerodynamic efficiency where traditional rotary- and fixed-wing designs suffer from severe Reynolds effects. Considerable progress has been achieved at the insect-scale (<<1 g) using high-bandwidth soft actuators such as piezoelectric bimorphs [2] and dielectric elastomer actuators [3], leading to milestone demonstrations of robotic flies and bees [4, 5, 6, 7, 8, 9]. However, these sub-gram platforms remain largely tethered or duration-limited due to the absence of power-dense materials capable of replicating insect biomechanics. To overcome this, researchers have designed small-scale FWMAVs under 100 g [1], integrating onboard computation and power within strict Size, Weight, and Power (SWaP) constraints. These systems, inspired by birds [10, 11, 12, 13, 14, 15, 16], bats [17, 18, 19, 20], beetles [21, 22, 23], dragonflies [24], fruit flies and more [25, 26, 27, 28, 29, 30, 31, 32, 33, 34], have significantly extended operational endurance and flight complexity, marking an important evolution in flapping-wing robotics.
@@
Amidst these advances, one of nature’s most aerodynamically sophisticated fliers, the butterfly, remains surprisingly underexplored. Butterflies employ a unique flight paradigm distinct from both fast-flapping insects and larger vertebrate flyers. Unlike birds or bats that rely on active wing morphing or tail actuation [35, 36, 28], or bees and fruit flies that flap above 150 Hz [37], butterflies generate sufficient lift and achieve controlled tailless flight using highly compliant wings driven anteromotorically by forewings and through large-amplitude strokes at only $\sim$10 Hz [38, 39]. Their characteristic low wing aspect ratios (<5<5) result in the highest wing area-to-body-mass ratios among flying animals [40]. Despite these seemingly unstable traits, butterflies exhibit extraordinary agility, such as executing rapid climbs, sharp turns, and erratic zigzags [41], suggesting a flight strategy where stability emerges from the dynamic coupling of compliant wings and body undulation rather than high flapping rate or auxiliary control surfaces.
@@
Existing aerodynamic studies have focused primarily on insects with rigid, high-aspect-ratio wings flapping at high frequencies. While butterflies also exploit unsteady mechanisms such as delayed stall [42], wake capture [43], and clap-and-fling [44], their biomechanical realization of these effects differs fundamentally. Their large and compliant wings interact strongly with airflow, leading to complex fluid-structure coupling and increased thoracic loading that ultimately lowers flapping frequency while amplifying unsteady effects. Despite the vast diversity of lepidopteran wing morphologies [45], the aerodynamic contributions of these traits remain poorly characterized [44]. The anteromotoric flight of butterflies further renders them inherently underactuated in six-degree-of-freedom (6-DOF), posing a long-standing challenge in understanding how tailless animals achieve maneuvers such as turning [46]. Moreover, their characteristic body undulation likely arises from intricate wing-body interactions that couple aerodynamic and inertial effects [47, 48, 49].
@@
Biological investigations have revealed complex flight behaviors such as climbing [50, 48], forward flight [51, 38], and turning [46, 52] via high-speed videography. Nonetheless, they remain mostly descriptive and lack quantitative links to controllable actuation. On the other side, Computational Fluid Dynamics (CFD) simulations can provide insight into flow structures [53, 54], but they often assume rigid wings and omit aeroelastic coupling and body-wing interaction, limiting their biological and engineering relevance. Due to the rudimentary understanding of biomechanics and the difficulties of modeling fluid-structure interactions and wing-body coupling [55, 56, 57], the translation of the unique mechanisms underlying butterfly flight into engineered systems has remained limited [58, 59, 60, 61, 54]. Butterfly-inspired, two-winged tailless FWMAVs—especially those with low-aspect-ratio wings and operating at low flapping frequencies—remain largely unexplored in autonomous robotics. This knowledge gap presents both a challenge and an opportunity to uncover the principles governing butterfly flight and translate them into controllable flapping-wing robots.
@@
In this work, we present AirPulse, a 26 g butterfly-inspired flapping-wing robot that, to the best of our knowledge, represents the first and lightest two-winged tailless FWMAV with fully onboard sensing and closed-loop control reported in peer-reviewed work (Fig. 1, A and B, Table 1, and movie S1). Despite the abundance of research on four-winged and tailed flapping-wing platforms, two-winged tailless systems with low-aspect-ratio wings remain poorly understood, primarily because their flight dynamics are dominated by strong fluid-structure coupling and the absence of tail-based stabilization. The AirPulse robot addresses this gap by replicating the fundamental aerodynamic and structural characteristics of butterflies, featuring low wing aspect ratio, biologically relevant wing loading—the lowest among reported studies, indicating efficient aerodynamics (Table 1)—and highly compliant wings reinforced by carbon fiber spars in a vein-inspired layout (Fig. 1G). This biomimetic reinforcement yields spatially heterogeneous stiffness, allowing compliant yet controllable wing deformation, thereby contrasting the uniform stiffness typical of rod-reinforced or rigid-wing designs [62, 63]. We also revealed that the large-amplitude flapping motion induces cyclic variations in the center of gravity and moment of inertia (Fig. 2, A and B), which manifest as pronounced body undulation, a main challenge for control in tailless flapping-wing flight.
@@
To govern this underactuated system, we systematically characterized the mapping between flapping wing modulation parameters—including amplitude, frequency, stroke angle offset, and stroke timing—and force-torque generation (Fig. 5). This quantitative mapping of control effectiveness fills a fundamental gap in understanding how such tailless flapping-wing robots allocate control authority in the absence of auxiliary surfaces, which prior studies have not explicitly addressed [64, 61, 54, 65, 66]. Furthermore, we introduced a novel rhythm generator, namely Stroke Timing Asymmetry Rhythm (STAR), which enables smooth and stable asymmetric wing motion, where the degree of asymmetry varies linearly with a single control parameter (Fig. 3, E and F). This formulation directly overcomes the discontinuity and instability issues found in prior stroke timing modulation approaches [65, 67, 68, 14], providing a mathematically grounded and physically robust foundation for precise stroke timing control.
@@
Building upon morphological prototype and rhythm modulation strategy, we developed an onboard control architecture that integrates Proportional-Integral-Derivative (PID)-based attitude regulation with real-time state estimation to cope with the inherent oscillatory dynamics of undulatory flight. In contrast to fixed-wing or bird-inspired robots that rely on steady aerodynamic moments, the AirPulse robot must continuously manage rapidly varying inertial and aerodynamic couplings throughout each wingbeat. Our controller was shown to maintain pitch and yaw stability even under large oscillatory motion of the body. Experimental validation demonstrated fully untethered, stable free flight using two flapping modulations (stroke angle offset and stroke timing based on STAR)—without tails or external control surfaces—achieving pitch and yaw tracking, climbing, and turning maneuvers with an average power consumption of $\sim$5.9 W and a power loading of 4.38 g/W (Fig. 6, D and E, and movie S4 and S5). These results showcase the first onboard controlled flight of a 26 g butterfly-inspired robot and establish a foundation for understanding and controlling tailless, underactuated flapping-wing systems with two low-aspect-ratio wings, thus paving the way for for non-invasive ecological monitoring and confined-space inspection where traditional rigid drones may fall short.
@@ Results
To investigate how butterfly-like wing kinematics and flapping modulation enable controlled free flight, we developed AirPulse (Fig. 1A), a small-scale flapping-wing robot inspired by the lime swallowtail (Papilio demoleus), one of the most widely distributed butterfly species worldwide. To the best of our knowledge, the AirPulse robot is the first and lightest butterfly-inspired flapping-wing platform reported to date with fully integrated onboard sensing and closed-loop control (flight-ready mass $\sim$26 g; dry mass $\sim$24 g; Fig. 1B and Table 1). It was designed as a biological counterpart, being lightweight, morphologically plausible, and instrumented for quantitative studies of aeroelastic wing-air interactions, flapping pattern generation, and feedback controller design.
@@
Within a 26 g total mass, the AirPulse robot features a biomimetic wing structure (wingspan $\sim$60 mm) connected to a compact body core (Fig. 1A and movie S1). Its lightweight design was achieved primarily through two complementary strategies. First, we adopted a novel body architecture in which all avionics are directly assembled on a single 3D-printed fuselage (Fig. 1C), unlike previous designs where the control unit or battery was mounted at the abdomen and connected to the thorax via a carbon rod [64, 38, 54, 66]. This integration eliminates structural redundancy and allows the Inertial Measurement Unit (IMU) to be co-located near the Center of Gravity (CG). Second, we minimized the wing surface area, yielding the smallest wing planform reported among butterfly-inspired robots in the literature (Table 1). Guided by scaling laws and geometric similarity [41], where $S\sim l^{2}$ and $m\sim l^{3}$ ($S$ is single wing area, $m$ is body mass, $l$ is characteristic length), we estimated the desired wing size based on the biologically observed mass ratio between wings and body. Specifically, we calculated the non-wing mass and divided it by 0.7, assuming the wings contribute approximately 30% of total mass in butterflies [59]. This provided an estimate of the total mass and required wing surface. Starting from a larger prototype capable of generating sufficient lift, we scaled the design accordingly and iteratively refined it through flight tests to identify the minimal wing area that maintained stable lift and control. This iterative process substantially reduced the overall mass while preserving flight capability. As a result, the wings of the AirPulse robot account for approximately 38% of the total body mass (Fig. 1B), comparable to that observed in dried specimens of Papilio xuthus [59], confirming the biological plausibility of the design.
@@
Each lateral wing assembly comprises a forewing and hindwing mechanically coupled via a 3D-printed connector to a common servo arm. Two micro servos mounted on the body core independently actuate the forewings (Fig. 1D), reproducing the anteromotoric wing actuation seen in real butterflies [69]. This configuration yields a coupled, synchronous forewing-hindwing motion with a characteristic phase lag (movie S2). For efficient actuation, the servo axes are canted relative to the longitudinal symmetry plane to align the torque transmission with the structural stiff axis. This configuration produces a fixed $50^{\circ}$ V-shaped dihedral, realized through the fuselage and servo mounts, which promotes inter-wing force asymmetry similar to that of real butterflies [54]. The resulting flapping motion passively induces body undulation with a phase offset between flapping and body pitch angle (Fig. 1D and Fig. 6, F and H). Due to the larger wing-to-body ratio of the robot, this oscillatory coupling is pronounced than in larger robotic fliers (bird-inspired drones), motivating control strategies that explicitly account for periodic body dynamics as presented later. Besides micro servos, the body core houses a custom flight control board featuring a 9-axis IMU and barometer for state estimation, direct Pulse Width Modulation (PWM) servo control, and per-rail current monitoring (Fig. 1C and Materials and Methods). All states are telemetered to the custom ground station for data logging. The IMU is positioned with the robot’s CG and aligned to the body frame (+x+x forward, +y+y left, +z+z up) to minimize inertial coupling and aerodynamic interference. A lightweight power management circuit supports both 1S and 2S LiPo batteries, allowing weight reduction with a trade-off in flight endurance.